OpenStudy (anonymous):

Use the Distributive Property and simplify the expression. 6 + 4(3x – 2) + 5 A. 12x + 9 B. 12x + 3 C. 7x + 9 D. 3x + 13

OpenStudy (anonymous):

B?? @200205650

OpenStudy (anonymous):

okay what are you getting when you distribute 4(3x – 2)

OpenStudy (anonymous):

4 to (3x-2) 6 + 4(3x-2) + 5 = 6 + 12x - 8 + 5 = [12x] + [6 - 8 +5] = 12x + 3 (B)

OpenStudy (anonymous):

Yes that is right! :)

OpenStudy (anonymous):

Correct !!

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

You're welcome!